# Comprehensive Assessment of Jailbreak Attacks Against LLMs

### v2

@@ Introduction
Figure 3: Fine-grained attack success rates for direct attacks of each method on various violation categories.
To delve deeper into the relationship between the unified policy, comprising the 16 violation categories, and the jailbreak attack taxonomy, which includes six distinct types of jailbreak attacks, we employ heatmaps to visualize the attack performance. The heatmaps for each model are available in Figure 3 and Figure 10 in Section A.6.
First, the heuristic-based, feedback-based, and parameter-based jailbreak attacks are relatively robust and versatile among all the violation categories. It is particularly noteworthy that LAA has achieved extremely high success rates (over 0.50) across all violation categories on all LLMs except Llama3.1. Among all the human-based jailbreak methods, AIM is the most effective one from various categories of our policy. The above conclusions intuitively support the findings we presented previously (see Section 6).
The heatmaps display a negative correlation, indicating that obfuscation-based jailbreak attacks (except DrAttack) consistently show poor performance across all violation categories in the policy. Our analysis of individual LLMs reveals that all obfuscation-based jailbreak attacks except DrAttack are effective only when applied to GPT-3.5 and GPT-4 across all violation categories. Even in the case of the most vulnerable LLM, Vicuna (see 10(a)), all obfuscation-based methods except DrAttack exhibit attack success rates lower than or equal to 0.50 across all violation categories, and in the vast majority of cases, even lower than 0.10. However, on GPT-4 (see 3(b)), the Zulu attack demonstrates remarkable effectiveness, with attack success rates exceeding 0.50 across all violation categories except for Privacy Breach. This reiterates our previous findings in Section 6.1 that the formidable capabilities of both GPT-3.5 and GPT-4 pose greater challenges for their alignment and safeguard mechanisms.
Takeaways
Our experimental results demonstrate that all the LLMs are vulnerable to jailbreak attacks. Heuristic-based and feedback-based attacks tend to be effective across all target LLMs, while most obfuscation-based attacks exhibit poorer generalization capabilities, being efficient only against specific models. For different violation categories from our unified policy and jailbreak taxonomy, the attack performances vary. In general, Vicuna is the most vulnerable LLM based on both our unified policy and jailbreak attack taxonomy. Among all attack methods, LAA performs the best, even only with black-box access. The human-based jailbreak method can still achieve a very high ASR score in the black-box scenario. Considering that many heuristic-based methods start with human-based jailbreak prompts, the human-based jailbreak methods should receive more attention. For the relationship between the policy and attack taxonomy, the heuristic-based, feedback-based, and parameter-based jailbreak attacks are more robust among all the violation categories. Additionally, we hope that these organizations can further optimize based on their proposed policies to mitigate jailbreak attacks better.
Ablation Study
@@
Commercial LLMs typically charge users based on the token counts used in their requests, and the token numbers significantly affect the LLMs’ response speed. As a result, adversaries may manage and optimize the token length of prompts to control costs when utilizing these models for jailbreaking. Figure 4 illustrates the average token numbers of jailbreak prompts used in different methods across six target models. The results of different models are available in Figure 11 in Section A.6.
The average token number of our baseline is the average token count of the forbidden questions, which is 14.78. Our results indicate that, for the black-box scenario, token counts of the human-based jailbreak prompt and many approaches that used this prompt as the initial prompt are significantly larger than others. For instance, the average token count of all human-based methods reaches surprisingly 676.79, and even the shortest one, AIM, also has an average token count of 382.78. Those methods using the human-based jailbreak prompt as the initial seed, including AutoDAN, GPTFuzz, LAA, and MasterKey, also need lots of tokens, with the average token counts all exceeding 300. However, feedback-based methods are not the case. PAIR and TAP have relatively short jailbreak prompts, as their initial seeds are not necessary to be those long jailbreak prompts in the wild. Meanwhile, GCG and COLD, which generate jailbreak prompts by adding fixed-length content, have the shortest prompt lengths among feedback-based methods. In contrast, human-based jailbreak approaches often adopt a more comprehensive strategy to circumvent LLM safeguards. These methods systematically examine a wide array of conditions and integrate them into the prompt. Techniques such as role-playing, reiterating the purpose, and specifying the output format are employed, resulting in prompts with large token numbers.
@@
Table 4: Time efficiency for different attacks on the full test dataset: Black-box attacks’ duration is averaged over eight LLMs, while others are averaged over five open-source LLMs.
Method Average Runtime (min) DrAttack 492 AutoDAN 699 GPTFuzz 338 LAA 513 GCG 2162 COLD 578 PAIR 701 TAP 772 AdvPrompter1 1324 Generation Exploitation 341
@@
    AdvPrompter’s running duration includes the time to fine-tune the prompter LLM and generate prompts. Generating 160 prompts costs about 40 minutes.
@@
As we know, most jailbreak attacks in human-based or obfuscation-based method only require a negligible amount of time for a content modification. These attacks can be launched swiftly as they have been collected as a continuously updated dataset [66]. Therefore, we treat their time consumption as zero. On the other hand, DrAttack, heuristic-based, feedback-based, fine-tuning-based and parameter-based jailbreak attacks typically demand more time and computational resources to conduct attacks. Therefore, it is important to consider the trade-off between attack effectiveness and time efficiency when evaluating jailbreak methods. We demonstrate the average time consumption of these methods in Table 4. Full results are available in Table 12 of Section A.6. Note that these results are preferred for qualitative analysis as many methods involve external API calls, influenced by uncontrollable factors like traffic limitations.

### v3

@@ Introduction
Figure 3: Fine-grained ASRs for direct attacks of each method on various violation categories (Llama3.1).
We employ heatmaps to visualize attack performance and analyze the relationship between the unified policy and the jailbreak attack taxonomy. The heatmaps for each LLM are available in Figure 3Figure 10 and Figure 11 in Appendix K.
The heatmaps show that the ASR trends on individual policies are generally consistent with the overall ASR trends. That is, taxonomies/methods with higher overall ASRs also tend to exhibit higher ASRs on individual policies. For example, LAA usually achieves the highest ASR across most policies, and obfuscation-based methods remain effective only on specific models.
However, we do observe some exceptions in certain models (e.g., LLaMA-3.1). For instance, in high-severity violation categories such as Terrorist Content, LAA’s ASR is even lower than that of obfuscation-based methods like Zulu. One possible reason is that such severe categories may involve super-enhanced safety guardrails targeting specific English phrases. Since LAA’s adversarial prompts always contain the original forbidden question, those sensitive phrases might still appear in the generated prompts, leading to significantly lower ASRs. In contrast, methods like Zulu or DrAttack either replace or obfuscate these specific phrases, which could contribute to their relatively higher ASRs in these cases.
We also find that, for strongly aligned LLMs like Llama3.1, on some vulnerable violation categories, the ASRs of jailbreak attacks are usually lower than the baseline, indicating that the jailbreak prefixes/suffixes themselves are also the target of such LLMs’ internal safeguards, which aligns with our discussion in Section 6.1.
Takeaways
First, in the most realistic black-box attack scenarios, jailbreak attacks can still pose substantial security threats to the latest models. Second, intra-category and inter-category attacks exhibit distinct patterns. Human-based methods play a crucial role, as they often serve as the source for initial seeds. Heuristic-based methods inherently depend on the initial seed, making them relatively non-robust. Feedback-based methods demonstrate better robustness. obfuscation-based attacks are often effective only against specific powerful LLMs. Last, strongly safety-aligned models could determine whether to reject user inputs based on both the question and the jailbreak prompt. This results in the weak robustness of human-based methods, as well as other approaches that rely on them (e.g., heuristic-based and fine-tuning-based attacks).
Ablation Studies
We systematically conduct comprehensive ablation studies, which include attack time efficiency, prompt token length, transferability, and attack performance on GPT-3.5 and GPT-4 over time. Our ablation studies reveal some hidden patterns, such as the heuristic-based attacks have good transferability, but their jailbreak prompts are relatively long. The details of ablation studies are in Appendix A.
Jailbreak Defenses
Table 4: Average ASRs across nine LLMs under different defenses (direct attack settings). “All” denotes that all eight defense methods are deployed together. The reduced values of ASRs compared with no additional defenses are recorded in the corresponding brackets. The highest value in each column is highlighted in bold blue.
Jailbreak
Method
SR Erase Moderation Perplexity PG LG LG2 LG3 All
AIM 0.54 (↓0.08) 0.03 (↓0.59) 0.61 (↓0.01) 0.62 (↓0.00) 0.00 (↓0.62) 0.32 (↓0.30) 0.36 (↓0.26) 0.33 (↓0.29) 0.00 (↓0.62)
Devmoderanti 0.38 (↓0.05) 0.04 (↓0.39) 0.42 (↓0.01) 0.43 (↓0.00) 0.00 (↓0.43) 0.13 (↓0.30) 0.21 (↓0.22) 0.06 (↓0.37) 0.00 (↓0.43)
Devmodev2 0.39 (↓0.04) 0.00 (↓0.43) 0.42 (↓0.01) 0.43 (↓0.00) 0.00 (↓0.43) 0.29 (↓0.14) 0.28 (↓0.15) 0.15 (↓0.28) 0.00 (↓0.43)
Base64 0.13 (↓0.03) 0.15 (↓0.01) 0.16 (↓0.00) 0.16 (↓0.00) 0.16 (↓0.00) 0.16 (↓0.00) 0.10 (↓0.06) 0.03 (↓0.13) 0.02 (↓0.14)
Combination 0.24 (↓0.04) 0.10 (↓0.18) 0.28 (↓0.00) 0.28 (↓0.00) 0.28 (↓0.00) 0.28 (↓0.00) 0.28 (↓0.00) 0.15 (↓0.13) 0.06 (↓0.22)
Zulu 0.25 (↓0.05) 0.30 (↓0.00) 0.30 (↓0.00) 0.04 (↓0.26) 0.29 (↓0.01) 0.30 (↓0.00) 0.29 (↓0.01) 0.24 (↓0.06) 0.04 (↓0.26)
DrAttack 0.55 (↓0.08) 0.58 (↓0.05) 0.63 (↓0.00) 0.63 (↓0.00) 0.57 (↓0.06) 0.59 (↓0.04) 0.59 (↓0.04) 0.41 (↓0.22) 0.36 (↓0.27)
AutoDAN 0.61 (↓0.09) 0.01 (↓0.69) 0.69 (↓0.01) 0.70 (↓0.00) 0.00 (↓0.70) 0.36 (↓0.34) 0.38 (↓0.32) 0.36 (↓0.34) 0.00 (↓0.70)
GPTFuzz 0.50 (↓0.08) 0.30 (↓0.28) 0.50 (↓0.08) 0.58 (↓0.00) 0.01 (↓0.57) 0.40 (↓0.18) 0.30 (↓0.28) 0.18 (↓0.40) 0.00 (↓0.58)
LAA 0.79 (↓0.08) 0.06 (↓0.81) 0.87 (↓0.00) 0.87 (↓0.00) 0.00 (↓0.87) 0.50 (↓0.37) 0.51 (↓0.36) 0.10 (↓0.77) 0.00 (↓0.87)
GCG 0.51 (↓0.06) 0.46 (↓0.11) 0.57 (↓0.00) 0.09 (↓0.48) 0.12 (↓0.45) 0.38 (↓0.19) 0.28 (↓0.29) 0.17 (↓0.40) 0.02 (↓0.55)
COLD 0.38 (↓0.07) 0.34 (↓0.11) 0.44 (↓0.01) 0.45 (↓0.00) 0.39 (↓0.06) 0.29 (↓0.16) 0.29 (↓0.16) 0.25 (↓0.20) 0.17 (↓0.28)
PAIR 0.57 (↓0.07) 0.33 (↓0.31) 0.63 (↓0.01) 0.64 (↓0.00) 0.56 (↓0.08) 0.46 (↓0.18) 0.37 (↓0.27) 0.33 (↓0.31) 0.16 (↓0.48)
TAP 0.59 (↓0.06) 0.35 (↓0.30) 0.65 (↓0.00) 0.65 (↓0.00) 0.59 (↓0.06) 0.50 (↓0.15) 0.43 (↓0.22) 0.38 (↓0.27) 0.19 (↓0.46)
Masterkey 0.50 (↓0.06) 0.00 (↓0.56) 0.56 (↓0.00) 0.56 (↓0.00) 0.00 (↓0.56) 0.27 (↓0.29) 0.29 (↓0.27) 0.27 (↓0.29) 0.00 (↓0.56)
AdvPrompter 0.26 (↓0.08) 0.24 (↓0.10) 0.34 (↓0.00) 0.34 (↓0.00) 0.29 (↓0.05) 0.18 (↓0.16) 0.13 (↓0.21) 0.12 (↓0.22) 0.04 (↓0.30)
Defense Methods
We widely test eight external defenses, including Self-Reminder (SR) [84], Moderation [47], Perplexity [3, 35], Erase [39], Llama-Guard (LG) [34], Llama-Guard-2 (LG2) [52], Llama-Guard-3 (LG3) [53], Prompt-Guard (PG) [54] (details in Appendix H).
Experiments
Metrics. We employ bypass rate (BR) and ASR as our evaluation metrics, the same as previous works [34, 39, 3, 35]. Other setups align with those in our main experiments.
Results. We report the average ASRs in Table 4 and BRs in Table 11 of Appendix E.2. First, none of the defenses can completely defend against all jailbreak attacks, as demonstrated by high BR and ASR in many cases. The lightweight Prompt-Guard model is extremely effective for human-based methods and all other approaches that utilize an initial seed. This includes all heuristic-based methods and MasterKey. Significantly, Prompt-Guard can lower the average ASR of nearly all these methods to zero. However, Prompt-Guard does not perform effectively against some other methods. For instance, even with Prompt-Guard active, the ASRs of DrAttack and TAP still reach 0.55 and 0.59, respectively. Moderation is almost ineffective in the majority of cases. Perplexity is effective on those jailbreak prompts with high perplexity (such as Zulu and GCG).
In addition, we compose all eight defense mechanisms together. The results show that all the human-based, heuristic-based, and other methods using initial seeds are almost ineffective, with ASRs close to zero, including the most powerful attack LAA. The reason is that jailbreak prompts in these methods are often derived from a fixed set of seeds, exhibiting similar patterns and distributions that differ from benign user inputs, making them easier to detect. DrAttack, COLD, PAIR, and TAP are still effective. These four methods do not rely on initial seeds and could craft more diverse and natural jailbreak prompts, making it more difficult for defense methods to capture prompt characteristics.
Discussion
Safety Alignment Trade-Offs. We notice that some violation categories (e.g., AI Usage Disclosure) have higher ASRs than others, even already covered in the providers’ usage policies. The baseline ASRs for such categories are also high. One reason may be that such categories seem to be “less harmful.” It is likely that during safety alignment (e.g., RLHF), human annotators paid less attention to these categories, leading LLMs to continue following instructions for them. The LLM providers may also make some trade-offs between utility and safety regarding these “less harmful” categories, despite their policies explicitly covering these categories. How to deal with such “less harmful” categories is still an open question.
Future Attacks and Defenses. Jailbreak prompts with natural and diverse patterns are harder to defend against, especially those that do not need initial seeds, which are stealthier and more resistant to defenses. In contrast, seed-based attacks are easily detected due to limited diversity. We hope the researchers, both the attack and defense sides, could prioritize attention on attacks requiring no initial seeds rather than focusing solely on the modification of known jailbreak prompts or their variants.
Conclusion
In this paper, we conduct a unified and comprehensive analysis of 17 representative jailbreak attacks and propose a novel attack taxonomy with six categories. We formulate a unified policy spanning 16 violation categories from five major LLM providers and build a diverse forbidden question dataset of 160 questions for experiments. Our ablation study highlights the unique features of each attack method beyond ASRs. Results show that under real-world black-box settings, the latest LLMs remain vulnerable to current jailbreak attacks, with LAA performing the best. Current defenses could effectively defend against those attacks using human-based initial seeds but struggle to defend against those not using such seeds. We call on the community to focus on creating and defending against jailbreak attacks that require no initial seeds, and hope our evaluation supports the development of trustworthy LLMs.
Limitation
Research Scope. According to popular research repositories [77, 95], there are now over 200 jailbreak attacks. It is infeasible to evaluate them all within a single paper. Although we try our best to include 17 representative attacks (see Section G.1) and uncover valuable patterns among the methods, we acknowledge that the research scope of the paper is still limited.
Static Policies and Questions. Previous harmful question datasets either rely on old policies or are based on authors’ self-proposed guidelines without supporting references. To fill the gap, we take the union of policies from multiple companies in 2024 to organize unified policies. Since not all models cover all policies, we encourage readers to use our results based on their use cases. We also acknowledge that our policies and corresponding datasets are static and may also become outdated as LLM-related policies evolve over time. We mainly analyze the inter-violation-category difference. However, we acknowledge that questions in the same category may also trigger different responses from LLMs. Investigating the intra-violation-category response difference, such as misinformation across different topics, deserves exploration in the future.
Jailbreak Evaluation Methods. Ideal evaluations of jailbreaking involve expert manual annotation, assessing both ASR and response quality. However, this approach is impractical due to high costs. We thus propose an automatic ASR evaluation method, which, while superior to others (see Section D.1), is still imperfect. Lacking domain knowledge, we cannot properly assess the quality of jailbroken responses or compare them with harmful knowledge from other sources. But we can confirm that LLM jailbreak methods significantly simplify the generation of harmful responses. Methods evaluating both ASR and response quality deserve more attention.
Potential Biases. Training of strong LLMs has almost exhausted all public data, and some data may inevitably have been used by newer models. Thus, we acknowledge that involving LLMs in building a forbidden dataset [96, 74] might introduce unknown biases, despite our manual checks and modifications. Additionally, many jailbreak attacks involve using other LLMs for assistance, such as ChatGPT, which could also introduce biases. Our human annotation may still introduce some unavoidable biases.
Ethical Considerations
In this study, we exclusively utilized data that is publicly accessible and did not engage with any participants. Therefore, it is not regarded as human subjects research by our Institutional Review Boards (IRB). However, our primary goal involves assessing the efficacy of various jailbreak methods, so we will inevitably reveal which methods can trigger inappropriate content from LLMs more effectively. Thus, we took great care to share our findings responsibly. We ensure that we will reveal our findings to the involved LLM service providers, including OpenAI, Google, ZhipuAI, LMSYS, DeepSeek AI, and Meta. In line with prior research [74, 82], we firmly believe that the societal advantages derived from our study significantly outweigh the relatively minor increased harm risks.
Acknowledgements
We thank all anonymous reviewers, ACs, and SACs for their constructive comments. This work is partially funded by the European Health and Digital Executive Agency (HADEA) within the project “Understanding the individual host response against Hepatitis D Virus to develop a personalized approach for the management of hepatitis D” (DSolve, grant agreement number 101057917) and the BMBF with the project “Repräsentative, synthetische Gesundheitsdaten mit starken Privatsphärengarantien” (PriSyn, 16KISAO29K).
Appendix A Ablation Studies
Transferability
Table 5: Average ASRs for transfer attacks. The baseline here refers to the average ASRs on the other eight LLMs (except Vicuna) without utilizing jailbreak techniques.
Method ChatGLM3 Llama2 Llama3 Llama3.1 GPT-3.5 GPT-4 DeepSeek-V3 PaLM2 Average
DrAttack 0.59 0.30 0.27 0.24 0.55 0.51 0.55 0.56 0.45
AutoDAN 0.87 0.39 0.30 0.29 0.58 0.34 0.80 0.82 0.55
GCG 0.39 0.33 0.27 0.29 0.44 0.36 0.45 0.27 0.35
COLD 0.35 0.30 0.28 0.28 0.40 0.28 0.45 0.20 0.32
GPTFuzz 0.76 0.13 0.19 0.23 0.41 0.45 0.75 0.36 0.41
LAA 0.82 0.21 0.36 0.35 0.99 0.71 0.85 0.75 0.63
PAIR 0.44 0.24 0.27 0.28 0.43 0.40 0.61 0.56 0.40
TAP 0.56 0.34 0.35 0.30 0.73 0.63 0.66 0.73 0.54
AdvPrompter 0.44 0.27 0.24 0.28 0.45 0.29 0.40 0.45 0.35
Average 0.58 0.28 0.28 0.28 0.55 0.44 0.61 0.52 0.44
Baseline 0.38 0.31 0.39 0.39 0.44 0.38 0.49 0.47 0.40
Table 6: Average ASRs of all jailbreak attacks (transfer attack) across different violation categories The baseline here refers to the average ASRs across different violation categories on the other eight LLMs (except Vicuna) without utilizing jailbreak techniques.
Violation Category ChatGLM3 Llama2 Llama3 Llama3.1 GPT-3.5 GPT-4 DeepSeek-V3 PaLM2 Average Baseline
Hate, Unfairness or Harassment 0.28 0.06 0.08 0.09 0.44 0.23 0.41 0.30 0.24 0.10
Malicious Software 0.43 0.10 0.00 0.01 0.31 0.26 0.53 0.54 0.27 0.10
Well-being Infringement 0.81 0.47 0.59 0.42 0.86 0.83 0.86 0.68 0.69 0.83
Physical Harm 0.36 0.03 0.00 0.00 0.36 0.22 0.46 0.39 0.23 0.10
Disinformation Spread 0.43 0.02 0.01 0.00 0.42 0.23 0.50 0.51 0.27 0.04
Privacy Breach 0.43 0.03 0.00 0.02 0.28 0.09 0.43 0.53 0.23 0.04
Adult Content 0.79 0.42 0.20 0.51 0.78 0.74 0.83 0.53 0.60 0.83
Political Activities 0.87 0.67 0.60 0.62 0.90 0.84 0.86 0.62 0.75 0.86
Impersonation 0.83 0.64 0.73 0.47 0.81 0.77 0.83 0.57 0.71 0.89
Terrorist Content 0.32 0.02 0.00 0.08 0.19 0.07 0.32 0.48 0.19 0.08
Unauthorized Practice 0.74 0.67 0.67 0.63 0.82 0.60 0.71 0.54 0.67 0.79
Safety Filter Bypass 0.57 0.19 0.26 0.28 0.57 0.41 0.59 0.47 0.42 0.28
Risky Government Decisions 0.52 0.04 0.10 0.28 0.37 0.25 0.54 0.63 0.34 0.30
AI Usage Disclosure 0.90 0.80 0.66 0.79 0.87 0.84 0.90 0.58 0.79 0.94
Third-party Rights Violation 0.52 0.27 0.52 0.22 0.59 0.43 0.62 0.48 0.46 0.29
Illegal Activities 0.44 0.00 0.06 0.11 0.33 0.24 0.40 0.51 0.26 0.03
(a) GPT-3.5
(b) GPT-4
(c) DeepSeek-V3
(d) PaLM2
Figure 4: Fine-grained ASRs for transfer attacks of each method on various violation categories (closed-source settings).
@@ Appendix A Ablation Studies
Figure 5: Fine-grained ASRs for transfer attacks of each method on various violation categories (open-source settings).
Evaluation on Attack Taxonomy. We first studied the attack transferability of different jailbreak methods. Table 5 demonstrates the transfer attack of different categories of our attack taxonomy on the rest of the LLMs. Surprisingly, we find that the attack performance of LAA drops minor on some LLMs. For example, it achieves ASRs over 0.70 on all LLMs except the Llama series. It could even achieve an ASR of 0.99 on GPT-3.5. For the other methods, the transferred jailbreak prompt is still effective on the rest of the models, but lower than the original attack performance. For instance, the average ASR score of AutoDAN is 0.55, higher than the baseline (0.40) but much lower than the original attack performance in Vicuna (0.98). In addition, for the white-box attacks, transferring the jailbreak attack can provide an effective solution against the LLMs with only black-box access. To illustrate, when jailbreaking PaLM2, AutoDAN demonstrates a notable ASR score of 0.82, meaning that this attack method exhibits good transferability on this model. GCG and COLD demonstrate relatively poor transferability, with average ASRs less than 0.35, even falling below the baseline.
Llama series models demonstrate robust resistance to transfer attacks, achieving average ASRs below 0.30, which falls even lower than the baseline, suggesting that they may have implemented tailored defenses against jailbreak prompts. In other words, this implies that Llama series models may not only detect harmful queries but also detect unusual characteristics associated with jailbreak prompts.
Evaluation on Unified Policy. We present the overall ASR results in Table 6 with different categories of the unified policy. In general, the transferred jailbreak prompts are still effective enough to launch the attacks. For instance, Political Activities still has a good average attack performance (0.75), similar to the original attack (0.78) in Vicuna. Notably, it can achieve a 0.90 ASR score to jailbreak GPT-3.5. The well-aligned Llama series models demonstrate strong resilience across most of the violation categories. Compared with the baseline, the average ASR of transfer attacks decreases across most violation categories.
Taxonomy-Policy Relationship. We also study the relationship between the unified policy and attack taxonomy under the transferability setting. We present the results for closed-source models in Figure 4. The results for open-sourced models could be found in Figure 5 in Appendix K.
We have observed that transfer attacks can boost the ASR across all challenging violation categories, including categories Illegal Activities, Privacy Breach, and Disinformation Spread, where the baseline ASRs are less than 0.05. Specifically, the average ASRs for transfer attacks in these categories have been increased to over 0.20.
Our detailed results for each model further elucidate the strong performance of AutoDAN, TAP, and LAA. As depicted in 4(a) and 4(b), transfer attacks conducted by AutoDAN, TAP, and LAA have improved ASRs compared to the baseline across most violation subcategories on GPT-3.5 and GPT-4, respectively. Note that transfer attacks have shown strong attack effectiveness on certain violation categories that could lead to serious consequences. For instance, TAP achieves an ASR success rate of 0.63 and 0.60 on Terrorist Content in GPT-3.5 and GPT-4, respectively. The high success rates of transfer attacks imply low-cost access to illicit resources or information, which is particularly concerning and warrants significant attention.
@@
Commercial LLMs typically charge users based on the token counts used in their requests, and the token numbers significantly affect the LLMs’ response speed. As a result, adversaries may manage and optimize the token length of prompts to control costs when utilizing these models for jailbreaking. Figure 6 illustrates the average number of tokens of jailbreak prompts used in different methods across six target models. The results of different models are available in Figure 12 in Appendix K.
The average token number of our baseline is the average token count of the forbidden questions, which is 14.78. Our results indicate that, for the black-box scenario, token counts of the human-based jailbreak prompt and many approaches that used this prompt as the initial prompt are significantly larger than others. For instance, the average token count of all human-based methods reaches more than 670, and even the shortest one, AIM, also has an average token count of 382.78. Those methods using the human-based jailbreak prompt as the initial seed, including AutoDAN, GPTFuzz, LAA, and MasterKey, also need lots of tokens, with the average token counts all exceeding 300. However, feedback-based methods are not the case. PAIR and TAP have relatively short jailbreak prompts, as their initial seeds do not necessarily need to be those long jailbreak prompts in the wild. Meanwhile, GCG and COLD, which generate jailbreak prompts by adding fixed-length content, have the shortest prompt lengths among feedback-based methods. In contrast, human-based jailbreak approaches often adopt a more comprehensive strategy to circumvent LLM safeguards. These methods systematically examine a wide array of conditions and integrate them into the prompt. Techniques such as role-playing, reiterating the purpose, and specifying the output format are employed, resulting in prompts with large token numbers.
@@
Table 7: Different methods’ runtime duration (minutes) for traversing the entire test dataset. These results are preferred for qualitative analysis as many methods involve external API calls, influenced by uncontrollable factors like traffic limitations. [^12] We do not consider the running time of DeepSeek-V3 as the API service is extremely unstable due to high workload and external attacks [18].
@@
As we know, most attacks in human-based or obfuscation-based methods only require a negligible amount of time for a content modification. These attacks can be launched swiftly as they have been collected as a continuously updated dataset [74]. Therefore, we treat their time consumption as zero. On the other hand, DrAttack, heuristic-based, feedback-based, fine-tuning-based, and generation-parameter-based jailbreak attacks typically demand more time and computational resources to conduct attacks. Therefore, it is important to consider the trade-off between attack effectiveness and time efficiency when evaluating jailbreak methods. We demonstrate the average time consumption of these methods in footnote 12. Note that these results are preferred for qualitative analysis, as many methods involve external API calls, influenced by uncontrollable factors like traffic limitations.